# I2VControl: Disentangled and Unified Video Motion Synthesis Control

### v2

@@ 3.3 Network, Training and Inference
During inference, we first use the Unidepth [26] metric depth estimation method to obtain the initial point set $\Omega$. Users then interact with our system to provide further input. They can first select the motion units of interest. Specifically, users can employ a bounding box and text description to select a small patch, and then set it as either a brush-unit or a drag-unit. If set as a drag-unit, users are required to provide a 6-DOF input to control the movement of that patch. Users can repeatedly select multiple patches until they have completed their design. Additionally, users can input an overall camera movement, which does not conflict with the motion of the selected patches. Our system provides a preview interface, allowing users to see the results of the drag and camera movement in real-time, achieving a “what you see is what you get” experience, which is highly convenient. After user interaction is complete, we convert the user inputs into control signals and conduct inference.
@@ 4 Experiments
Figure 4: Comparison on trajectory control task.
In this section, we introduce our experiments. Sec. 4.1 presents our implementation details and settings. In Sec. 4.2, we conduct the feature comprehensiveness comparison with previous methods to show the completeness and superiority of our framework design. In Sec. 4.3, we show our results and experimental comparisons. We strongly recommend visiting our project page to view video results: https://wanquanf.github.io/I2VControl.
@@ 4.1 Settings
Implementation Details. In this work, we utilize an Image-to-Video variant of the Magicvideo-V2 [35] as our foundational model ($24$ frames with resolution $704\times 448$). Training is executed on 16 NVIDIA A100 GPUs, with each GPU processing a batch size of $1$. The entire training spans over $100$K steps. Throughout the training phase, we maintain the parameters of the foundational model unchanged while focusing solely on the training of our newly incorporated adapter structure. For comparisons, we test on both motion strength based (motion brush) and motion trajectory based (motion drag, camera movement) tasks.
Datasets. We collect a training set of $970$K video clips, where each video may contain object motion or camera movement. To test on the motion trajectory based control task, we follow previous work DragAnything [38] and test on the VIPSeg [24] dataset (We selected all samples with at least $24$ frames in the testing set of VIPSeg, a total of $101$ testing samples). Given that the VIPSeg videos generally lack aesthetic appeal, we think them inadequate for real-world user scenarios. To address this, we invite several real-world users familiar with video creation to construct a manual labelled dataset, which should be more closely aligned with typical user scenarios. Specifically, they employ text-to-image model to generate some high-aesthetic images, and then manually construct the segmentation mask, the camera movement and object motion as the control signal, which completely simulates user operation. To test on the motion trajectory control task, we collect $30$ labelled samples with movable objects. Similarly, to test on the motion strength based control task, we construct a dataset of $63$ testing samples, $33$ of which are object motion (e.g., humans, animals) and the other $30$ are fluid effects (e.g., smoke, fire).
Metrics. For the motion trajectory control task, we refer to DragAnything [38] and TrackGo [13], evaluating the quality of the results by ObjMC. Note that since we allow control via dense trajectories instead of sparse control points, we compute this metric on dense trajectories. To evaluate motion brush control task, we firstly compute a motion strength score, denoted as MSC, as in I2VControl-Camera. Then we refer to SAM [20] to employ a IoU metric to evaluate whether the generated motion follows the user input mask. Moreover, for all experiments, we additionally compute the FID score to measure the quality of the results.
@@ 4.2 Feature Comprehensiveness Analysis
In this section, we theoretically compare the feature comprehensiveness between our method and some very recent baseline methods in Tab. 1. We mainly focus on the following control strategies: motion drag, camera extrinsics, and motion brush, all of which are important components of popluar video generation commercial products, such as Gen2 [28]. For motion drag, users can select some points/objects on the input image and drag them to target positions, thus guide the video generation process. For camera extrinsics control, users are allowed to set a camera pose sequence to decide the camera movement of the generated video. For motion brush, we adopt the definition from motion-I2V [29], which enables the selected region to move (without specifying a trajectory or direction), while not affecting other unselected areas. From Tab. 1, we can see that only our method can deal with all the above common control situations in one single framework.
It is noteworthy that although some motion drag methods (e.g. DragAnything [38], MOFA-Video [25]) do not accept a camera matrix as input, they propose that the effect of camera movement can be achieved by applying 2D drags to multiple points in the background. This also reminds us that both camera movement and motion drag can essentially be considered as 2D trajectory based control, where the former acts on the background and the latter on the subject. In the following sections, we conduct comparisons in two major categories: trajectory based control and motion brush.
@@ 4.3 Results and Comparisons
In this section, we present our experimental results and comparisons with previous methods. Secondly, in Sec. 4.3.1, we show the diverse capabilities of our framework by applying it on several inspiring and creative control situations. In order to form a more convincing experiment, we compare with other methods on trajectory based control in Sec. 4.3.2 and motion brush in Sec. 4.3.3.
@@ 4.3.1 Diverse Capabilities and Creative Exploration
As an disentangled and unified framework, I2VControl can combine multiple spatial control ideas without conflict, which can greatly meet the needs of video creation users. We show some examples in Fig. 3. The first three columns represent several situations of 6-DOF object movement, including basic translation and rotation. Note that the second column is a translation along the Z-axis, and the third column is the rotation of the head, which are difficult to achieve in previous methods. The fourth column shows an arbitrary camera movement. The fifth and sixth columns are motion brushes for objects and fluids, respectively. In the seventh column, we show a combined effect, where the camera and object movements, and the motion strength based brush are employed in one single sample. The eighth column shows a creative control from user idea. Specifically, we keep the foreground still and the background zoomed out, making it look like the background is quickly moving backwards (the Hitchcockian camera movement). Due to constraints on time and page space, we did not show more inspiring ideas. However, the examples provided sufficiently demonstrate that our system empowers users with the tools they need to unleash their creativity, encouraging them to explore and devise more intriguing control combinations.
4.3.2 Comparison on Trajectory Control
We compare on the trajectory control based task (including object motion and camera movement) with very recent methods DragAnything [38] and MOFA-Video [25]. As described in Sec. 4.1, we test on the VIPSeg [24] dataset and another manual dataset. For each sample, we divide the motion units on the initial frame. To construct the control signal for the comparing methods, we extract representative trajectories on each single units, where one trajectory is selected for the drag-units and four trajectories are selected for the borderland part to represent the camera movement. Quantitatively, we calculate the metric between the generated results and the ground truth control signals, and show the results on both datasets in Tab. 2. Compared to previous methods, we achieve better results on both VIPSeg and the manual dataset and both the ObjMC and FID metrics, which shows that our method is superior to previous methods on both controlling precision and fidelity. The qualitative results can be seen in Fig. 4. In the left sample, we use a pan-right camera movement and drag the object to the right, and only our result is right, thanks to our dense and disentangled nature. The sample on the right shows our good performance for rotation control.
@@ 4.3.2 Comparison on Trajectory Control
Table 2: Comparison on the trajectory control task.
4.3.3 Comparison on Motion Brush
Figure 5: Comparison on motion brush for fluid effects.
We compare on the motion brush control task, where we brush masks on images and set scalar motion strength for the selected region without a motion trajectory or direction guide. As described in Sec. 4.1, we test on two types of inputs, the movable objects and fluid effects. We outperform previous method Motion-I2V on both datasets and all metrics (see Tab. 3), reflecting our effectiveness better response to masks and more significant motion. The numerical difference is very intuitively reflected in the visualization results, see Fig. 6 and Fig. 5. The result of Motion-I2V has a small motion range, and the movable pixels seems not able to move out of the given mask. This might be related to the control stragety based on optical flow. In contrast, our method can produce very large-scale motion. For example, in Fig. 5, the flame emitted from the wand can be transformed into different shapes. The brushed objects can also move away from the origin mask, such as the animal on the right of Fig. 6.
Datasets Object Movements Fluid Effects
Methods Motion-I2V Ours Motion-I2V Ours
@@ 4.3.3 Comparison on Motion Brush
Table 3: Comparison on the motion brush control task.
@@ 5 Conclusion
In this work, we explore the controllability of video synthesis, a vital aspect for enhancing end-user utility. Our proposed framework, I2VControl, implements flexible video motion controllers with disentangled control signals, achieving a user-friendly interface and outstanding performance in various motion control tasks including motion strength, motion brushing, object dragging, and camera pose controlling. For the future, we aim to enhance the framework to handle increasingly complex scenarios, integrating more nuanced control over video motion not only at the structural but also at the textual level. Additionally, we also expect ongoing advancements in model architectures to boost the capabilities and operational efficiency of our framework.
Acknowledgement
We thank Yuxi Xiao and Yudong Guo for their generous help and advice with the tracking algorithm.

### v3

@@ 5 Experiments
Figure 8: Comparison on dragging control task.
In this section, we introduce our experiments. Sec. 5.1 presents our implementation details and settings. Sec. 5.2 presents the feature comprehensiveness comparison. In Sec. 5.3, we show our results and experimental comparisons. Please visit our anonymous page for video results.
@@ 5.1 Settings
Implementation Details. In this work, we utilize an Image-to-Video variant of the Magicvideo-V2 [37] as our foundational model (242424 frames, resolution 704×448704$\times$ 448704 × 448).
Datasets. We train on a training set of 970970970K video clips. To ensure rigorous comparison, we conduct comprehensive evaluation on 666 datasets. To maintain consistency with prior research, we adopt 333 widely-used existing datasets. Additionally, to provide a richer and more convincing comparison, we also curate 333 new datasets to address certain gaps in the functionalities of previous benchmarks. We invite several real-world users familiar with video creation to construct high-quality manual control signals including segmentation mask, camera movement and object motion, completely simulating user operation. The testing sets are summarized as:
  • VIPSeg [25]. Following previous works [40, 14], we test on VIPSeg (all testing samples with at least 242424 frames).
  • RealEstate10K [51]. Following previous works [38, 15, 8], we use the testing set of RealEstate10K.
  • MO [8]. Following previous work [8], we use MO dataset, containing rich camera movement and motions.
  • Manual Dragging. We invite real-world users to label samples of camera movement and object motion.
  • Object Brush. We invite users to label motion masks on input images of object motion (e.g., humans, animals).
  • Fluid Brush. We invite users to label motion masks on input images of fluid effect samples (e.g., smoke, fire).
Metrics. For a rigorous comparison, we strictly adopt metrics from previous works. To judge camera control effect, we follow previous works [38, 15, 8], using RotErr and TransErr. To judge dragging, we refer to previous works [40, 14], using ObjMC. To judge motion brush, we firstly refer to previous work [8] to compute a motion strength score MSC, and then refer to SAM [21] to employ a IoU metric to evaluate whether the generated motion follows the user input mask. Moreover, we additionally employ FID and a User study score to measure the quality.
@@ 5.2 Feature Comprehensiveness Analysis
In this section, we theoretically compare the feature comprehensiveness between our method and some very recent baseline methods in Tab. 1. For motion drag, users can select some points/objects on the input image and drag them to target positions, thus guide the video generation process. For camera extrinsics control, users are allowed to set a camera pose sequence to decide the camera movement of the generated video. For motion brush, we adopt the definition from motion-I2V [30], which enables the selected region to move (without specifying a trajectory or direction), while not affecting other unselected areas. From Tab. 1, we can see that only our method can deal with all the above common control situations in one single framework.
@@ 5.3 Results and Comparisons
In this section, we present our experimental results and comparisons. In Sec. 5.3.1, we show the diverse capabilities of our framework. Then we compare with previous methods in Sec. 5.3.2. We also provide our results on another base model to demonstrate our adaptive nature in Sec. 5.3.3.
@@ 5.3.1 Diverse Capabilities and Creative Exploration
As an disentangled and unified framework, I2VControl can combine multiple spatial control ideas without conflict, which can greatly meet the needs of video creation users. We show some examples in Fig. 7 (please see the captions in the figure for textual explanation). The examples provided sufficiently demonstrate that our system empowers users with tools to unleash their creativity, encouraging them to explore and devise more intriguing control combinations.
Figure 9: Comparison on camera control.
5.3.2 Comparison with Previous Works
We compare with recent methods including DragAnything [40], MOFA-Video [26], MotionCtrl [38], CameraCtrl [15], and Motion-I2V [30]. It is difficult to compare all methods under exactly the same experimental setup (the base models are inconsistent; even if the base models are aligned, the training data are not the same). We strive to reduce inconsistencies and train MotionCtrl and CameraCtrl on the same base model and data as ours. For the remaining methods, although the base models are different, they are at a similar level, all of which are U-Net methods (rather than recently developed high-quality DiT models). Furthermore, we will analyze the effect differences and point out that these differences are obviously not caused by the base model, but by the methods themselves.
Due to methodological limitations of the comparing methods, it is infeasible to compare them with ours under a unified control (like the 777-th sample of Fig. 7), which is exactly the best proof of our superiority. Nevertheless, we still choose suitable datasets for them to enable meaningful (albeit suboptimal) comparisons with us.
$\bullet$ VIPSeg and Manual Dragging Datasets. We compare with very recent methods DragAnything [40] and MOFA-Video [26] on the VIPSeg [25] dataset and Manual Dragging (short as MD) dataset. For each sample, we divide the motion units on the initial frame. To construct the control signal for the comparing methods, we extract representative trajectories on each single units, where one trajectory is selected for the drag-units and four trajectories are selected for the borderland part. Quantitative results are in Tab. 2, which shows that our method is superior to previous methods on both controlling precision and fidelity. The qualitative results can be seen in Fig. 8. In the left sample, we use a pan-right camera movement and drag the object to the right, and only our result is right, thanks to our dense and disentangled nature. The sample on the right shows our good performance for rotation control.
@@ 5.3.2 Comparison with Previous Works
[…]
Table 2: Comparison on VIPSeg and MD datasets.
$\bullet$ RealEstate10K and MO Datasets. We compare with MotionCtrl and CameraCtrl, on RealEstate10k (short as RE10K) and MO datasets in Tab. 3. Our quantitative results are the best, proving the precision and quality of our results. The qualitative results are shown in Fig. 9, showing our good control precision again.
Dataset Metric MotionCtrl CameraCtrl Ours
@@
[…]
Table 3: Comparison on RealEstate10K and MO datasets.
Figure 10: Comparison on motion brush for fluid effects.
$\bullet$ Object/Fluid Brush Datasets. We compare on the motion brush control task, where we brush masks on images and set scalar motion strength for the selected region without a motion trajectory or direction guide. We test on two datasets, the movable objects and fluid effects. We outperform previous method Motion-I2V on both datasets and all metrics (see Tab. 4), reflecting our effectiveness better response to masks and more significant motion. The numerical difference is very intuitively reflected in the visualization results, see Fig. 11 and Fig. 10. The result of Motion-I2V has a small motion range, and the movable pixels seems not able to move out of the given mask. This might be related to the control stragety based on optical flow. In contrast, our method can produce very large-scale motion. For example, in Fig. 10, the flame emitted from the wand can be transformed into different shapes. The brushed objects can also move away from the origin mask, such as the animal on the right of Fig. 11.
Dataset Metric Motion-I2V Ours
[…]
Table 4: Comparison on Object/Fluid Brush datasets.
@@
5.3.3 Adaptive Nature
To show our adaptive nature, we also train on other base models. Specifically, we train on Seedance [9](see visual results here) and Wanx 2.1 [32](we have released our code base on Wan 2.1). Please note that the implementation of attention varies depending on the specific base model used. However, the organization of the control signals and the encoder remain consistent. Our structure can still work well on these base models, which proves our adaptive nature.
@@ 6 Conclusion
In this work, we explore the motion controllability of image-to-video synthesis, a vital aspect for enhancing end-user utility and solving logic conflicts in previous methods. Our proposed framework, I2VControl, implements flexible video motion controllers with disentangled control signals, achieving a user-friendly interface and outstanding performance in various motion control tasks including motion brush, object dragging, and camera pose controlling. For the future, we aim to enhance the framework to handle increasingly complex scenarios, integrating more nuanced control over video motion not only at the structural but also at the textual level. Additionally, we also expect ongoing advancements in model architectures to boost the capabilities and operational efficiency of our framework.
Acknowledgement We thank Yuxi Xiao and Yudong Guo for their generous help and advice with the tracking algorithm. We also thank Hongrui Cai for developing the open-source version of the model and training code. Finally, we thank Jianzhu Guo for insightful suggestions on the gradio demo.